find three consecutive integers such that the sum of the first and the third integer is 8 less than the second integer.
Answer by ptaylor(2198)

Let x=the first
Then x+1=the second
And x+2=the third
Sum of the first and third is:
x+x+2 or 2x+2
Eight less than the second is:
(x+1)-8 or x-7
Now we are told that the sum of the first and third is eight less than the second. So:
2x+2=x-7 subtract x and also 2 from both sides
2x-x+2-2=x-x-7-2 collect like terms
x=-9--------------first
x+1=-9+1=-8----------second
x+2=-9+2=-7----------------third
CK
the sum of the first and third is eight less than the second:
2x+2=x-7
2*(-9)+2=-9-7
-16=-16

Hope this helps---ptaylor